OpenStudy (anonymous):

Whitney left school and drove 7 miles to meet her mother at a restaurant. Her mother left at the same time and drove 8 miles, traveling 5 mi/h faster than her daughter. If the two arrived at the restaurant at the same time, what was Whitney’s speed?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (zpupster):

w = the speed of Whitney m = the speed of her mother m = w + 5 7/w = 8/m cross multiply 7m = 8w

OpenStudy (zpupster):

plug m = w + 5 into the last equation 7(w + 5) = 8w

OpenStudy (anonymous):

ok

OpenStudy (zpupster):

7w+35 = 8w 7w-7w+35 =8w-7w w=35